# Multi-head committees enable direct uncertainty prediction for atomistic foundation models

### v1

@@ Abstract
Machine learning potentials have become a standard tool for atomistic materials modelling. While models continue to become more generalisable, an open challenge relates to efficient uncertainty predictions for active learning and robust error analysis. In this work, we utilise MACE and its multi-head mechanism to implement a committee neural network potential for message-passing architectures, where the committee comprises multiple output modules attached to the same atomic environment descriptors. As with traditional committees of independent networks, the standard deviation of the predictions functions as an estimate of the model’s uncertainty. We show for a range of datasets in custom-build models that the uncertainty of the force predictions correlates well with the true errors. We subsequently apply this concept to foundation models, specifically MACE-MP-0, where we train only the newly attached output heads while keeping the remaining part of the model fixed. We use this approach in an active learning workflow to condense the training set of the foundation model to just 5% of its original size. The foundation model multi-head committee trained on the condensed training set enables reliable uncertainty estimation without any substantial decrease in prediction accuracy.
@@ I Introduction
While proving to be a major leap in the field for the exploration of new material chemistry and physics, foundation models are in most applications only qualitatively correct and can still show unphysical behaviour [Focassio2025/10.1021/acsami.4c03815]. In this context, it would be advantageous to have easy ways of quantifying a foundation model’s uncertainty, which could be used to assess and monitor the accuracy of its predictions and identify failures more directly [Imbalzano2021/10.1063/5.0036522, Dai2025/10.1515/revce-2024-0028]. A range of methods and workflows to address this issue is known in the context of MLPs [Gawlikowski2023/10.1007/s10462-023-10562-9], but they have not yet seen widespread adoption in foundation models, given the extensive cost of training. Established methods for uncertainty prediction in MLPs include a last-layer approximation of prediction rigidities [Bigi2024/10.1088/2632-2153/ad805f], the prediction of confidence intervals using quantile regression [Bilbrey2025/10.1038/s41524-025-01572-y], Gaussian mixture models trained on atomic environment descriptors [Zhu2023/10.1063/5.0136574], a model-free estimator based on information entropy [Schwalbe-Koda2025/10.1038/s41467-025-59232-0], committee neural network [Schran2020/10.1063/5.0016004, Carrete2023/10.1063/5.0146905] potentials, and shallow ensembles [Kellner2024/10.1088/2632-2153/ad594a]. Once available, uncertainties can be used for active learning, a data-driven workflow to find the most relevant training structures out of a large set of candidates [Schran2020/10.1063/5.0016004, Carrete2023/10.1063/5.0146905, Schaaf2023/10.1038/s41524-023-01124-2, Holzmuller2023/activeLearning]. Such uncertainty estimates can then be used to monitor the reliability of a model’s prediction or, in the context of active learning, to condense and optimise training sets and to reduce the number of necessary reference electronic structure calculations. Having this uncertainty for foundation models would be particularly valuable. Considering the enormous effort required to train a foundation model, a solution to the uncertainty quantification problem should take advantage of the capacity of the foundation model while leaving its predictive power untouched. Furthermore, it should not add a substantial computational cost to the model and allow for a calculation of the uncertainty on the fly.
@@
II Results
In order to enable simple uncertainty prediction within the MACE framework, we have implemented an MHC architecture, shown schematically in panel a) of Figure 1. The general idea of such a committee has previously been outlined by Kellner et al. [Kellner2024/10.1088/2632-2153/ad594a], and a rudimentary implementation of it has been tested for MACE by Bilbrey et al. [Bilbrey2025/10.1038/s41524-025-01572-y]. While details of this architecture are discussed in full in the Methods Section, we give here a short, high-level summary of the main concepts. Leveraging the existing multi-head functionality within MACE, initially conceptualised in the context of fine-tuning [Batatia2023/10.48550/arXiv.2401.00096], we train multiple output heads to different subsets of the total training set. Two options for distributing the training data are illustrated in panel b) of Figure 1. One evenly distributes the full training set between the output heads (“disjoint”), while the other picks the subsets randomly and independently of each other from the full training set (“overlapping”). Having multiple readouts enables us to use the standard deviation between head predictions for uncertainty quantification, similar to committee neural network potentials, while requiring little additional architectural overhead. In addition, this design choice makes it very easy to extend a packaged foundation model without requiring retraining. For the upcoming results, it should be noted that in every calculation, the model’s hyperparameters can influence the results in many unintended, subtle ways, for example the shape of the correlation distribution. We have kept the hyperparameters as consistent as possible between models to minimise these factors.
Below, we demonstrate the power of this methodology by applying it to a series of systems and models of increasing complexity. We start with a gas-phase molecule, move to the condensed phase with liquid water, explore chemical space with a model for multiple organic molecules, and finally enhance a foundation model with uncertainty prediction.
@@ II Results
3BPA
As a first step, we show that the MHC can be used to estimate the uncertainty of MLP predictions. To this end, we test and compare the two different options for subsampling the MHC training set. We then compare these against the baseline of a naive committee, comprising individual full MACE models each trained on randomly chosen subsets of the training data. For these first tests, we use the 3BPA (3-(benzyloxy)pyridin-2-amine) dataset [Kovacs2021/10.1021/acs.jctc.1c00647], which contains structures of the drug-like 3BPA molecule from MD trajectories at different temperatures. In the 3BPA molecule, shown in the top left panel of Figure 2, the most important degrees of freedom are the torsions along the bonds connecting the two six-membered rings. The training data is based exclusively on $300\text{\,}\mathrm{K}$ structures, whereas the test data originates from sampling at $1200\text{\,}\mathrm{K}$ and therefore reaches a broader region of configuration space than the training set. Figure 2 shows the correlation between the actual error in force or energy per atom and the corresponding standard deviation of the committee predictions. The distribution of the values is shown as a two-dimensional histogram using hexagonal bins on a logarithmic scale. In order to map to the true generalisation error, these standard deviations were scaled in postprocessing using a scaling factor calculated from the validation set [Imbalzano2021/10.1063/5.0036522], as detailed in the Methods section. The individual scaling factors $\alpha$ are given in the upper left corner of every panel of Figure 2. To better illustrate the overall trend, we also bin the distribution along the uncertainty axis to contain an equal number of data points in each bin and calculate the mean error for each bin, shown as orange lines. A reference gray line indicates perfect correlation between errors and uncertainty. We see a significant spread of the distribution surrounding the ideal line, while the binned averages follow the optimal correlation closely. For all models investigated, there are data points with high uncertainty but small error, which is not an issue, as a model can produce an accurate prediction “by accident”, despite high uncertainty. Importantly, there are no instances where the model’s error is high while the uncertainty is low, which would be a clear indication of an unreliable uncertainty estimator. For all of our committee models, there is a clearly empty zone in the bottom right corner, especially for forces, which means that an increased error will always be indicated by an increased uncertainty.
Figure 2: The correlation between the RMSEs of the models and the scaled committee disagreements for forces per atom and energy of the whole molecule. The scaling parameter $\alpha$ is given in the top left corner of each subplot. The orange line indicates the binned average RMSE for all atoms or structures for forces and energies, respectively, binned along the committee disagreement. All bins contain the same number of data points and are, therefore, not of equal size. The gray line shows perfect correlation between uncertainty measure and actual error. The 3-(benzyloxy)pyridin-2-amine (3BPA) molecule used in this analysis is shown in the inset of the first plot.
Finally, we use the 3BPA system to analyse the prediction accuracy of the different model types. The evolution of the force and energy RMSEs shown in Figure 3 illustrates that the naive committee consistently outperforms the MHCs in prediction accuracy. However, this is expected, as the naive committee has almost 8 times as many trainable parameters as the MHCs due to the full independence of every committee member. Comparing the two different versions of MHCs shows a slight advantage in terms of accuracy for the one with overlapping training sets. In this version, each multi-head training set consists of 80% of the training data for each head for the overlapping committee and just 1/8 of the set for the disjoint committee. The training set for each head comprises 80% of the training data for the overlapping committee and just 12.5% of the training data for the disjoint committee. We conclude that the smaller size of the dataset available to each head in the disjoint committee leads to this small discrepancy. In the bottom row of Figure 3, the difference between the error of the whole committee and the average error of the committee members is shown. For the naive committee, the committee consistently outperforms the individual members due to the increased number of trainable parameters. For the overlapping committee, the differences are small and independent of the number of training structures. The improvement in performance typically associated with committee MLPs remains absent, as the MHC is conceptually similar to adding a dropout layer to the output blocks. While classic dropout layers set the output of random nodes to zero during training, the MHC does so for all nodes connected to certain output heads. For every structure in the training set, the same nodes are consistently removed in every epoch of the training process. It would be unreasonable to expect a significant performance boost from combining multiple of such predictions. Therefore, no significant improvement in predictive power should be expected. Contrary to this expectation, the disjoint model shows a noticeable improvement when introducing the committee. This improvement stems from the number of training examples each output head sees during training. Especially for small training datasets, the subsets for each head are extremely sparse, negatively affecting the prediction accuracy. With a large overall training set size, this sparsity decreases, leading also to more similar single and committee predictions.
Liquid water
Figure 4: Correlation of committee force disagreement and RMSE for each atom in a 64-molecule box of liquid water. The top row shows results for structures with classical nuclei — the same distribution as the training data. The bottom row shows results for structures with quantum nuclei — a distribution different from the training data.
rMD17
While each of the previous test cases focused on a single molecular system, in this section, we expand our investigation to cover a more diverse set of molecules within one model to probe uncertainty predictions across chemical compound space. To test this, we employ the rMD17 [Christensen2020/10.1088/2632-2153/abba6f] dataset comprising MD structures sampled at $500\text{\,}\mathrm{K}$ of 10 different organic molecules (consisting of H, C, N, and O atoms), which are shown in Figure 5. We randomly selected 50 structures from each of the 10 molecules to form our training set and 1000 structures of each molecule for separate per-species test sets. The unscaled mean and standard deviation of the uncertainties are shown in Figure 5 against the force errors for the multi-head and naive committees for all 10 molecules individually, to show how the model can handle predictions of different complexities. For an easier comparison, we plot the binned averages of all three approaches in one subplot. The shaded area indicates the standard deviation of the data in each bin to illustrate the spread of the data. The magnitudes of the uncertainty are dependent on the molecule under investigation, but the general trends are very consistent, as are the comparisons between the committee types. Analogously to our previous findings, all three approaches exhibit a similar correlation when plotting the mean force error for structures within a certain range of disagreement, and only the scale of these curves differs. As expected, the naive committee shows the highest level of uncertainty due to the large differences between its members, and the overlapping MHC shows the lowest, as its members are the least diverse. We also conducted a correlation analysis using the Pearson correlation coefficient and relative log-likelihood [Kellner2024/10.1088/2632-2153/ad594a], coming to the same conclusions (details in the SI, Section S2). Unfortunately, the correlation between energy uncertainty and energy errors is substantially weaker, as discussed in further detail in the SI S1. We also use the rMD17 dataset to examine committee disagreement for unknown atomic systems, as this will be especially relevant in the context of foundation models. Therefore, we remove one molecule from the training data and train the committees on the remaining 9 molecules. When we examine the committee disagreement on a test set of the molecule that was taken out of the training data, we find that the uncertainty still correlates well with the errors but can be distinctively low for certain atoms. We attribute this to the local environment of these atoms, which is similar to the local environment of atoms present in the training data. We discuss these results in more detail in the SI S3. Importantly, we find that our previous conclusions are also valid for heterogeneous datasets and that committee disagreement works well with large, diverse training sets and out-of-domain test structures.
Figure 5: Correlation between the unscaled committee disagreement and the force RMSE for individual atoms in the systems. The curves are plotted individually for all 10 different molecules in the rMD17 dataset and the structure of the respective molecule is shown above the plots. The curves show the average force RMSE binned along the committee disagreement, as the orange curves in Figure 2 and 4. The shaded areas around the lines are the standard deviations within each bin.
MACE MP-0 foundation model
After showing that the MHC provides a reliable measure of prediction uncertainty, we adapt it to a foundation model. We start from the MACE-MP-0b foundation model [Batatia2023/10.48550/arXiv.2401.00096] trained on the MPtrj dataset [Deng2023/10.1038/s42256-023-00716-3], which contains 146k crystalline structures calculated using density functional theory at the PBE+U [Perdew1996/10.1103/PhysRevLett.77.3865] level. We then equip this foundation model with an MHC by adding eight new output heads with a random initialisation of weights to the existing pre-trained head. As with custom-build models, the committee prediction is the mean of all new output heads, excluding the original head. When training the MHC output heads, we leave the weights of the AED block and the original head fixed. To obtain a more compact training set for the MHC, we use an iterative QbC workflow based on its committee disagreement [Schran2020/10.1063/5.0016004]. This reduces the original dataset to 8,000 structures — just 5% of its original size. This condensed training dataset is divided across the eight committee heads using the “disjoint” distribution, resulting in 1,000 training structures for each head. By using a much smaller training set and taking advantage of the pre-trained model, the computational cost of training is greatly reduced. Overall, we are adding roughly 15,000 parameters to the model, accounting for less than 0.2% of the total model size. It took 37 hours on an NVIDIA Hopper H100 GPU to execute the 800,000 optimisation steps of the whole MHC, a small fraction of the 2,600 GPU hours of the original model (trained on NVIDIA A100 GPUs across multiple nodes) [Batatia2023/10.48550/arXiv.2401.00096]. Especially for the smaller foundation models, this training effort can also be performed on consumer-grade GPUs.
Figure 6: Correlation between the committee disagreement and the force error for the modified MACE MP-0b model, where the output heads were trained on a reduced sample of the MPtrj dataset containing 8000 structures. The first subplot is on test data from the MPtrj dataset, which is not included in the reduced dataset. For the two subsequent plots, 10,000 structures from the NVT and NpT dataset of the OMAT database were used. As in the previous figures, the orange lines indicate the average force RMSE, binned along the direction of the uncertainty measure.
For evaluation, we used 10,000 out-of-distribution structures taken from the NVT and NpT OMAT datasets [Barroso-Luque/10.48550/arXiv.2410.12771], which are calculated with the same reference method as the MPtrj dataset, and 10,000 structures from the MPtrj dataset not selected during QbC. Figure 6 shows the correlation between the actual force errors per atom and the uncertainty prediction for the three datasets in question. Compared to previous cases, the results are more spread out, as both the training and test data are much more diverse. In the case of OMAT, this means that many of the systems in the test set are not present in the training data at all. The MPtrj dataset, on the other hand, functions as an indicator for the in-data regime, as it is not independent of the training data of the model. The AEDs of the model were trained on the full MPtrj dataset, including the structures of this test set. Overall, the committee disagreement correlates well with the error of the committee prediction and therefore functions as a reliable uncertainty measure even for foundation models. However, unlike with custom-made models, we observe occasional instances where a low uncertainty coincides with a high error. This is most likely due to structures in the test set, which have similar characteristics to some structures in the training set, resulting in uniform predictions by the committee. In the MPtrj dataset, where the structures of the test set were used to train the original foundation model, but not the committee output heads, this effect is particularly pronounced. However, overall, these instances are isolated enough not to deteriorate the reliability of the MHC uncertainty prediction for foundation models.
An important question that remains is whether the new committee model predictions are still accurate compared to the original, despite being trained on a small fraction of the original dataset. To test whether QbC is advantageous compared to other, computationally cheaper options, we created two alternative training sets. During the QbC run to select the reduced training dataset, it is noticeable that structures with high force components are selected at a much higher rate than structures close to the equilibrium. Therefore, we created a training set of the 8,000 structures with the highest force per atom. Additionally, we also randomly selected 8,000 structures from the initial dataset to form a third training set. Both datasets were used to train MHC models in the same way as the QbC dataset. Figure 7 shows the distribution of the force errors for the original MACE-MP-0 model and the new MHC models. As expected, the original MACE-MP-0 model has the highest prediction accuracy of all models, but the advantage over the re-trained models is small. Among the models trained on less data, the model with the QbC selected data performs the best, but the differences compared to other models are modest. It should be noted that for the OMAT datasets, a very small number of predictions had an extremely high error for all models, including the original MACE-MP-0 model. Due to the nature of root mean square errors, these few predictions dominate the final RMSE. Therefore, we decided to exclude all force components for which the error of the original MACE-MP-0 model is higher than $5\text{\,}\mathrm{e}\mathrm{V}\mathrm{/}\AA$ . This amounts to a total of 764 force components, or 0.019% of the combined OMAT test sets. For the MPtrj dataset, the original MACE-MP-0 model has a much stronger advantage over the retrained models than for the OMAT data. This is because the OMAT data is entirely unseen by all models, whereas the MPtrj test set was part of the MACE-MP-0 training dataset, and therefore, this strong performance should be expected. This also explains why the randomly selected subset outperforms the QbC selected set, as we sampled it from the structures not picked during QbC. Therefore, some structures of the randomly selected training set are included in this test set. As we show in the SI section S4, condensing the training set even further will eventually lead to a measurable decrease in prediction accuracy.
Overall, we conclude that our approach of adding an MHC to a foundation model preserves its predictive power. This is, on one hand, achieved by leaving the original foundation model — including its output head — intact. On the other hand, we have shown that the added output heads, trained on a condensed training dataset, also largely retain the foundation model’s performance, as most of the predictive capacity lies within its AEDs, which were extensively trained on the full MPtrj dataset. Optimising the weights of the AEDs further on a condensed training set would degrade the model’s performance.
III Discussion
The second part of this work used the MHC in active learning to condense a reduced training set out of the original MPtrj dataset, a widely used training set for foundation models. We selected 8000 out of 146k structures from the MPtrj dataset and used them to train the output heads of an adapted medium-sized MACE-MP-Ob foundation model. The AEDs of the model were left untouched to preserve its expressive capabilities. We showed that the MHC based on foundation models displays a good correlation between committee disagreement and actual force error, even though instances where the uncertainty is underestimated are slightly more common. When comparing the predictions of the original MACE-MP-0b model with the new model, we saw only a small drop in performance, although the output heads were trained on only 5.5% of the full training dataset. We also examined the criteria for selecting the reduced training sets and found that the QbC-selected model performs the best, even though the advantage is moderate. Given the compact nature of the condensed data sets, this opens up the possibility to obtain foundation models trained on expensive high-level electronic structure methods by recalculating the previously condensed training set and optimising only the output blocks. QbC can help ensure that the performance stays as close as possible to the original foundation model. Overall, the uncertainty estimation of the MHC architecture introduced here further increases the robustness of foundation models.
IV Methods
Multi-head committee
@@ IV Methods
Distribution of training data
To increase the heterogeneity between the multiple heads, we train each head on a different subset of the complete training set. We consider two possible strategies: either randomly sampling a fraction of the total set of structures for every head, or evenly distributing the whole dataset across the heads without any overlap between the subsets. Both strategies are illustrated in panel b) of Figure 1 and bring different benefits. The first strategy, in which a fraction of the total set of structures is randomly sampled for each head, is called “overlapping” due to the overlaps between resulting training sets. Its potential downside is that in the final concatenated training set, on which the common parts of the model are trained, structures appear multiple times without an even distribution among them. Therefore, the AEDs will be trained more often on some structures than others. The second strategy, in which the whole dataset is evenly distributed among the heads, is called “disjoint” as the training subsets do not overlap. As shown in Figure 3, there can be situations where the datasets used to train the output heads of the model are too sparse, and the prediction accuracy consequently decreases significantly if the original dataset is already extremely small. The main benefit of the disjoint strategy is that it would increase the diversity between the output heads more strongly, mitigating known problems of a common bias between the committee members [Kahle2022/10.1103/PhysRevE.105.015311].

### v2

@@ Abstract
Machine learning potentials have become a standard tool for atomistic materials modeling. While models continue to become more generalizable, an open challenge relates to efficient uncertainty predictions for active learning and robust error analysis. In this work, we utilize MACE and its multi-head mechanism to implement a committee neural network potential for message-passing architectures, where the committee comprises multiple output modules attached to the same atomic environment descriptors. As with traditional committees of independent networks, the standard deviation of the predictions functions as an estimate of the model’s uncertainty. We show for a range of datasets in custom-build models that the uncertainty of the force predictions correlates well with the true errors. We subsequently apply this concept to foundation models, specifically MACE-MP-0, where we train only the newly attached output heads while keeping the remaining part of the model fixed. We use this approach in an active learning workflow to condense the training set of the foundation model to just 5% of its original size. The foundation model multi-head committee trained on the condensed training set enables reliable uncertainty estimation without any substantial decrease in prediction accuracy.
@@ I Introduction
While proving to be a major leap in the field for the exploration of new material chemistry and physics, foundation models are in most applications only qualitatively correct and can still show unphysical behavior [Focassio2025/10.1021/acsami.4c03815]. In this context, it would be advantageous to have easy ways of quantifying a foundation model’s uncertainty, which could be used to assess and monitor the accuracy of its predictions and identify failures more directly [Imbalzano2021/10.1063/5.0036522, Dai2025/10.1515/revce-2024-0028]. A range of methods and workflows to address this issue is known in the context of MLPs [Gawlikowski2023/10.1007/s10462-023-10562-9], but they have not yet seen widespread adoption in foundation models, given the extensive cost of training. Established methods for uncertainty prediction in MLPs include a last-layer approximation of prediction rigidities [Bigi2024/10.1088/2632-2153/ad805f], the prediction of confidence intervals using quantile regression [Bilbrey2025/10.1038/s41524-025-01572-y], Gaussian mixture models trained on atomic environment descriptors [Zhu2023/10.1063/5.0136574], a model-free estimator based on information entropy [Schwalbe-Koda2025/10.1038/s41467-025-59232-0], committee neural network [Schran2020/10.1063/5.0016004, Carrete2023/10.1063/5.0146905] potentials, and shallow ensembles [Kellner2024/10.1088/2632-2153/ad594a]. Once available, uncertainties can be used for active learning, a data-driven workflow to find the most relevant training structures out of a large set of candidates [Schran2020/10.1063/5.0016004, Carrete2023/10.1063/5.0146905, Schaaf2023/10.1038/s41524-023-01124-2, Holzmuller2023/activeLearning]. Such uncertainty estimates can then be used to monitor the reliability of a model’s prediction or, in the context of active learning, to condense and optimize training sets and to reduce the number of necessary reference electronic structure calculations. Having this uncertainty for foundation models would be particularly valuable. Considering the enormous effort required to train a foundation model, a solution to the uncertainty quantification problem should take advantage of the capacity of the foundation model while leaving its predictive power untouched. Furthermore, it should not add a substantial computational cost to the model and allow for a calculation of the uncertainty on the fly.
@@
II Methods
In order to enable simple uncertainty prediction within the MACE framework, we have implemented an MHC architecture. The general idea of such a committee has previously been outlined by Kellner et al. [Kellner2024/10.1088/2632-2153/ad594a], and was explored for MACE foundation models with an emphasis on energies by Bilbrey et al. [Bilbrey2025/10.1038/s41524-025-01572-y]. Leveraging the existing multi-head functionality within MACE, initially conceptualized in the context of fine-tuning [Batatia2023/10.48550/arXiv.2401.00096], we train multiple output heads to different subsets of the total training set. We lay out two options for distributing the training data: One evenly distributes the full training set between the output heads (“disjoint”), while the other picks the subsets randomly and independently of each other from the full training set (“overlapping”). Having multiple readouts enables us to use the standard deviation between head predictions for uncertainty quantification, similar to committee neural network potentials, while requiring little additional architectural overhead. In addition, this design choice makes it very easy to extend a packaged foundation model without requiring retraining.
Multi-head committee
@@ II Methods
Distribution of training data
During training, each structure gets assigned to a specific head and is subsequently used to optimize the AED and the assigned head. If we want to use a structure for multiple heads, it will appear multiple times in the training set with different assigned heads. To increase the heterogeneity between the heads, we train each of them on a different subset of the complete training set. We consider two possible strategies: either randomly sampling a fraction of the total set of structures for every head, or evenly distributing the whole dataset across the heads without any overlap between the subsets. Both strategies are illustrated in panel b) of Figure 1 and bring different benefits. The first strategy, in which a fraction of the total set of structures is randomly sampled for each head, is called “overlapping” due to the overlaps between resulting training sets. Its potential downside is that in the final concatenated training set, on which the common parts of the model are trained, structures appear multiple times without an even distribution among them. Therefore, the AEDs will be trained more often on some structures than others. The second strategy, in which the whole dataset is evenly distributed among the heads, is called “disjoint” as the training subsets do not overlap. As will be shown in the Results section, there can be situations where the datasets used to train the output heads of the model are too sparse, and the prediction accuracy consequently decreases significantly if the original dataset is already extremely small. The main benefit of the disjoint strategy is that it increases the diversity between the output heads more strongly, mitigating known problems of a common bias between the committee members [Kahle2022/10.1103/PhysRevE.105.015311].
@@
Pearson correlation coefficient
The Pearson correlation coefficient [Pearson1895/10.1098/rspl.1895.0041]
$r(\epsilon,\sigma)=\frac{\sum_{A}\left(\epsilon(A)-\overline{\epsilon}\right)\left(\sigma(A)-\overline{\sigma}\right)}{\sqrt{\sum_{A}\left(\epsilon(A)-\overline{\epsilon}\right)^{2}}\sqrt{\sum_{A}\left(\sigma(A)-\overline{\sigma}\right)^{2}}}$ (6)
is a measure of correlation between two datasets, in our instance, the error $\epsilon$ and the uncertainty $\sigma$. In the case of energy predictions, this uses the energy error and uncertainty per structure, and in the case of force predictions, it uses the root mean square error and mean uncertainty of all three force components of one atom. $\overline{x}$ denotes the mean over the whole dataset. The value of the correlation coefficient can range from -1 to 1, corresponding to perfect anti-correlation and perfect correlation, respectively, and 0 meaning no correlation at all.